OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 (5/3,4/3) (-5/3,-4/3) (-3/5,-3/4) (3/4,3/5)

OpenStudy (michele_laino):

Hint: I substitute the value of y from the second equation into the first one, so I get: \[x + (2x + 2) = - 3\]

OpenStudy (michele_laino):

what is x?

OpenStudy (anonymous):

idk?? D:D

OpenStudy (anonymous):

D:

OpenStudy (michele_laino):

hint: I can simplify that above equation as follows: \[\begin{gathered} x + 2x + 2 = - 3 \hfill \\ 3x + 2 = - 3 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

D: :(

OpenStudy (michele_laino):

\[3x = - 5\]

OpenStudy (michele_laino):

Now I divide both sides by 3, and I get: \[x = - \frac{5}{3}\]

OpenStudy (omarbirjas):

I think its B

OpenStudy (anonymous):

thxx <3

OpenStudy (michele_laino):

please wait, we have to compute y

OpenStudy (omarbirjas):

Dont trust my judgement listen to michele

OpenStudy (michele_laino):

hint: \[y = 2 \times \left( { - \frac{5}{3}} \right) + 2 = ...?\]